# MAGIK: Mapping to Analogous Goals via Imagination-enabled Knowledge Transfer

### v1

@@ 3.1 Reinforcement Learning
RL enables the agent to make decisions by interacting with an environment modelled as a Markov Decision Process (MDP). An MDP is defined by a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P$ is the transition dynamics, $r$ is the reward function and $\gamma\in[0,1)$ is the discount factor. Upon each interaction, the agent observes a state $s\in\mathcal{S}$, selects an action $a\in\mathcal{A}$, receives a reward $r(s,a)$, and transitions to a new state $s^{\prime}\sim P(\cdot\mid s,a)$. The agent’s goal is to learn a policy $\pi(a\mid s)$ that maximises the expected cumulative discounted return $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$.
@@ 4 Problem Setup
In addition, we make an assumption about the structure of the observations: each observation $x\in\mathcal{D}$ can be factorised into two mutually exclusive and exhaustive latent variables—a continuous variable $z$ encoding task-agnostic features (e.g., background, layout), and a discrete variable $c$ encoding task-specific information (e.g., goal-related object identity). Formally, we define the decomposition as $x=z\cup c$ with $z\cap c=\varnothing$. This assumption enables structured representation learning, allowing the agent to selectively manipulate task-relevant features while preserving the invariant structure across tasks.
Each task is modeled as a Markov Decision Process (MDP), defined as $\langle S,A,R,T,\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $R:S\times A\to\mathbb{R}$ is the reward function, $T:S\times A\to S$ is the transition function, and $\gamma\in[0,1]$ is the discount factor. Let $\pi_{\theta}(a|s)$ be a policy parameterized by $\theta$, which selects action $a\in A$ from state $s\in S^{\mathcal{S}}$ in the source task (denoted by superscript $\mathcal{S}$). Our goal is to infer an analogous policy $\pi_{\theta}(a|\tilde{s})$ for the target task, where $\tilde{s}\in S^{\mathcal{T}}$ represents an observation in the target domain.
To enable this, we learn an imagination network $\psi(\tilde{s})$ such that $\psi(\tilde{s})=s^{\prime}\in S^{\mathcal{S}}$—a state aligned with the source domain but conditioned on the target task. This allows the agent to apply the source policy as if it were operating in the source environment. Thus, MAGIK introduces a generative mapping module $\psi$ that transforms observations from the target domain into the source domain conditioned on the target task, enabling the agent to reuse its learned policy and induce behavior suitable for the new task.
@@ 5 Methodology
Formally, let $\pi_{\theta}(a|s)$ be the policy learned in the source task. The analogous policy in the target task is obtained by a VAE parameterized by $\psi$, which maps the target observation $\tilde{s}$ to its source-aligned counterpart. This allows the agent to execute the policy as $\pi_{\theta}(a|\psi(\tilde{s}))$.
Since the VAE operates in a semi-supervised manner, we assume that a small portion of the dataset is labeled by a human. These labeled samples play a crucial role in guiding the VAE to structure its latent space meaningfully. Labels correspond to different transferable elements that define the task (e.g., apple vs. orange). Additional details on the labeling are provided in Section 6.1.
@@ 5.1 Training of the Imagination Network
The VAE is modified to incorporate two latent variables: $z$ and $c$. The variable $z$ is a continuous latent code sampled from a Gaussian distribution and is designed to capture task-agnostic features, such as background structure or spatial layout. The variable $c$, on the other hand, is a discrete latent code representing task-specific factors, such as object identity or class-relevant information, and is sampled using the Gumbel-Softmax trick [18, 22]. These two variables are enforced to capture mutually exclusive and exhaustive aspects of the observation, formalised by the decomposition $x=z\cup c$ with $z\cap c=\varnothing$.
During reconstruction, the decoder combines the task-agnostic latent $z$ with the task-specific label $c$ inferred from the input observation to reconstruct the original image.
The VAE is trained by maximising the evidence lower bound (ELBO). As discussed earlier, our VAE employs two latent variables to disentangle class-agnostic and class-related information. The ELBO is derived by ensuring that $z$ captures only class-independent aspects of the observation while $c$ focuses solely on class-dependent features. To guide this latent separation, we utilise a small set of labelled observations. The overall ELBO consists of terms for both labelled and unlabeled data, with the final ELBO being their sum.
For labeled data, the ELBO is given by:
$\log p_{\theta}(x,y)\geq\underbrace{\mathbb{E}_{q_{\phi}(z|x,y)}[\log p_{\theta}(x|z,y)]}_{\text{Reconstruction term}}+\underbrace{\mathbb{E}_{q_{\phi}(c|x)}[\log p_{\theta}(y)]}_{\text{Supervision term}}\\ -\underbrace{KL(q_{\phi}(z|x)\parallel p_{\theta}(z))-KL(q_{\phi}(c|x)\parallel p_{\theta}(c))}_{\text{KL regularisation term}}.$ (3)
For unlabeled data, where supervision is not available, the ELBO simplifies to:
@@
where $K_{z}$ and $K_{c}$ are the centered kernel matrices corresponding to $z$ and $c$, respectively. Tr represent the trace of the matrix and $n$ is the number of samples. A lower HSIC value indicates reduced statistical dependence between $z$ and $c$, promoting better disentanglement. (The experimental evaluation is done by traversing through both the latent and results provided in the supplementary materials  B.)
@@ 5.2 Imagination Network Architecture
4:$z^{\text{target}}\sim q_{\psi}(z\mid x^{\text{target}})$triangleright Extract task-agnostic latent
@@
6:$\hat{x}^{\text{source-aligned}}\sim p_{\phi}(x\mid z^{\text{target}},c^{\text{source}})$triangleright Reconstruct with source task identity
@@
8:$a^{\text{target}}\sim\pi_{\text{source}}(\hat{x}^{\text{source-aligned}})$
@@ 6.1.1 Evaluating H1: Transferability
In the Reacher environment, the source task requires the agent to reach a blue target location, while the target tasks involve reaching other coloured targets—specifically red, green, and yellow. To enable transfer, the imagination network modifies observations from the target tasks by replacing the target colour with blue, effectively allowing the agent to interpret the new task through the lens of the source task. This substitution enables the learned policy to transfer without additional fine-tuning. The results of this transfer are summarised in Table $3$However, we found a performance drop if the source training data is not diverse enough. Details are provided in the Supplementary Section G.
@@ Appendix A Derivation of ELBO
As discussed in Section 5.1, we derive two ELBOs: one for the labelled dataset and another for the unlabelled dataset. We first derive the ELBO for the unlabelled case, and a similar approach follows for the labelled case with minor modifications.
A.1 Labelled Case
In the labelled case, we aim to maximize $\log p(x,y,z,c)$:
$\log p(x,y,z,c)$ $=\log\int_{z}\sum_{c}p(x,y,z,c)\,dz$ (7)
$\geq\int_{z}\sum_{c}\log p(x,y,z,c)\,dz.$ (8)
Since the integral over $z$ is intractable, we apply variational inference:
$\log p(x,y,z,c)$ $\geq\int_{z}\sum_{c}\log\left(q(z,c|x)\frac{p(x,y,z,c)}{q(z,c|x)}\right)dz$ (9)
$=\int_{z}\sum_{c}\log\left(q(z,c|x)\frac{p(x|y,z,c)p(y|z,c)p(z|c)p(c)}{q(z,c|x)}\right)dz.$ (10)
To enforce independence assumptions, we impose the following constraints:
  • $q(z,c|x)=q(c|x)q(z|x)$, assuming conditional independence between $z$ and $c$.
  • $p(x|y,z,c)=p(x|z,c)$, assuming $y$ does not influence $x$ given $z,c$.
  • $p(y|z,c)=p(y)$, since $y$ is obtained solely from human feedback.
  • $p(z|c)=p(z)$, assuming $z$ is independent of $c$.
Substituting these into the above expression and applying Jensen’s inequality, we obtain:
$\log p(x,y,z,c)$ $\geq\mathbb{E}_{q(z|x)q(c|x)}\left[\log p(x|z,c)\right]+\mathbb{E}_{q(c|x)}\left[\log p(y)\right]$ (11)
$\quad-KL(q(z|x)\parallel p(z))-KL(q(c|x)\parallel p(c)).$ (12)
A.2 Unlabelled Case
In the unlabelled case, we maximize $\log p(x,z,c)$, where:
$p(x,z,c)=p(x|z,c)p(z)p(c).$ (13)
Following a similar derivation as the labelled case, we obtain:
$\log p(x,z,c)\geq\mathbb{E}_{q(z|x)q(c|x)}\left[\log p(x|z,c)\right]-KL(q(z|x)\parallel p(z))-KL(q(c|x)\parallel p(c)).$ (14)
@@ Appendix D Training hyper-parameters
Reconstruction loss weight $2$
label loss weight $5$
Gradient clipping (Max norm) $1$
@@ Appendix F Imagination using GAN
Here, $z_{\text{reversed}}=\text{GRL}(E(x^{(l)}))$, and the gradient is scaled by $-\lambda_{\text{dis}}$. The total Encoder-Decoder loss is:

### v3

@@ 3.1 Reinforcement Learning
RL enables the agent to make decisions by interacting with an environment modelled as a Markov Decision Process (MDP). An MDP is defined by a tuple $\langle\mathcal{S},\mathcal{A},T,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:S\times A\to\mathbb{R}$ is the reward function, $T:S\times A\to S$ is the transition function, and $\gamma\in[0,1]$ is the discount factor. Upon each interaction, the agent observes a state $s\in\mathcal{S}$, selects an action $a\in\mathcal{A}$, receives a reward $R(s,a)$, and transitions to a new state $s^{\prime}\sim P(\cdot\mid s,a)$. The agent’s goal is to learn a policy $\pi(a\mid s)$ that maximises the expected cumulative discounted return $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$.
@@ 4 Problem Setup
In addition, we assume that each observation $x\in\mathcal{D}$ can be explained by two latent variables that capture complementary factors of variation: a continuous variable $z$, encoding task-agnostic features (e.g., background, layout), and a discrete variable $c$, encoding task-specific information (e.g., goal-related object identity). These variables are designed to be mutually exclusive (each accounts for a distinct aspect of the observation) and collectively exhaustive (together they are sufficient to reconstruct $x$). This assumption enables structured representation learning, allowing the agent to selectively manipulate task-relevant features while preserving invariant structure across tasks.
[…]
@@ 5 Methodology
Formally, let $\pi_{\theta}(a|s)$ be the policy learned in the source task. The analogous policy in the target task is obtained by a VAE parameterized by $\psi$, which maps the target observation $\tilde{s}$ to its source-aligned counterpart. This allows the agent to execute the policy as $)\pi_{\theta}(a|\psi(\tilde{s}$)).
Since the VAE operates in a semi-supervised manner, we assume that a small portion of the dataset is labeled by a human. We demonstrated, the labeled samples are randomly selected from the dataset. However, to further reduce the burden of human annotation, an active sampling strategy could be employed, where only the most informative or uncertain samples as determined by the classification head are queried for human labels. This would ensure that human supervision is concentrated on the most ambiguous cases, thereby improving label efficiency and scalability of the approach. These labeled samples play a crucial role in guiding the VAE to structure its latent space meaningfully. Labels correspond to different transferable elements that define the task (e.g., apple vs. orange). Additional details on the labeling are provided in Section 6.1.
@@ 5.1 Training of the Imagination Network
The Variational Autoencoder (VAE) is modified to incorporate two latent variables: $z$, a continuous latent code sampled from a Gaussian distribution to capture task-agnostic features (e.g., background structure or spatial layout), and $c$, a discrete latent code sampled using the Gumbel-Softmax trick [18, 22] to represent task-specific factors (e.g., object identity or class-relevant information). These variables are designed to capture mutually exclusive and exhaustive aspects of the observation, with their disentanglement enforced using a small set of labelled observations and the Hilbert-Schmidt Independence Criterion (HSIC) as an auxiliary loss term.
During reconstruction, the decoder combines $z$ and $c$, inferred from the input observation, to reconstruct the original image.
The VAE is trained by maximizing the evidence lower bound (ELBO), which balances reconstruction accuracy and regularization for both labelled and unlabelled data. The overall objective sums the ELBO terms for these data types, enabling the model to learn robust, disentangled representations across supervised and unsupervised settings.
For labelled data, the ELBO is given by:
$\log p_{\theta}(x,y)\geq\underbrace{\mathbb{E}_{q_{\phi}(z|x,y)}[\log p_{\theta}(x|z,y)]}_{\text{Reconstruction term}}+\underbrace{\mathbb{E}_{q_{\phi}(c|x)}[\log p(y|c)]}_{\text{Supervision term}}\\ -\underbrace{KL(q_{\phi}(z|x)\parallel p(z))-KL(q_{\phi}(c|x)\parallel p(c))}_{\text{KL regularisation term}}.$ (3)
Here, the supervision term is crucial: since we typically define $p(y\mid c)$ as a deterministic one-hot distribution, any probability mass assigned by $q_{\phi}(c|x)$ to an incorrect class yields a large penalty in the ELBO. Equivalently, this reduces to maximising $\log q_{\phi}(c=y\mid x)$, i.e., the log-likelihood of the correct class under the variational posterior. This makes the supervision term identical to the standard cross-entropy loss, ensuring that the discrete latent $c$ is aligned with semantic class identity rather than arbitrary clusters. We refer the reader to the supplementary Section A for detailed explanation.
For unlabelled data, where supervision is not available, the ELBO simplifies to:
@@
where $K_{z}$ and $K_{c}$ are the centered kernel matrices corresponding to $z$ and $c$, respectively. Here, Tr denotes the matrix trace and $n$ is the number of samples. A lower HSIC value indicates reduced statistical dependence between $z$ and $c$, promoting better disentanglement. (The experimental evaluation is done by traversing both latent variables; results are provided in the supplementary materials, Section B.)
@@ 5.2 Imagination Network Architecture
[…]
@@
[…]
@@
[…]
@@ 6.1.1 Evaluating H1: Transferability
In the Reacher environment, the source task requires the agent to reach a blue target location, while the target tasks involve reaching other coloured targets—specifically red, green, and yellow. To enable transfer, the imagination network modifies observations from the target tasks by replacing the target colour with blue, effectively allowing the agent to interpret the new task through the lens of the source task. This substitution enables the learned policy to transfer without additional fine-tuning. The results of this transfer are summarised in Table 3333However, we found a performance drop if the source training data is not diverse enough. Details are provided in the Supplementary Section G.
@@ Appendix A Derivation of ELBO
As discussed in Section 5.1, each observation is assumed to be explained by two complementary latent variables: a continuous latent $z$ that captures task-agnostic factors and a discrete latent $c\in\{1,\dots,K\}$ that captures task-specific factors. Together, these variables are designed to be mutually exclusive (each explains a distinct source of variation) and collectively sufficient to reconstruct the observation.
We now formalise this intuition within a probabilistic framework. The generative model for $x$ is parameterized by $\bm{\theta}$, while the variational distributions are parameterized by $\bm{\phi}$. The priors and the label model $p(y\mid c)$ are fixed and unparameterized. Formally, the joint model is given by:
$p(x,y,z,c)=p_{\bm{\theta}}(x\mid z,c)\,p(y\mid c)\,p(z)\,p(c),$
with the variational approximation factorised as:
$q_{\bm{\phi}}(z,c\mid x)=q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x).$
For completeness, a step-by-step derivation of the ELBO for both labelled and unlabelled cases is provided below.
A.1 Labelled Case ($x,y$ known)
We aim to maximize the marginal likelihood:
$\log p(x,y)=\log\int_{z}\sum_{c=1}^{K}p_{\bm{\theta}}(x\mid z,c)\,p(y\mid c)\,p(z)\,p(c)\,dz.$
We introduce the variational approximation $q_{\bm{\phi}}(z,c\mid x)=q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x)$ and apply Jensen’s inequality:
$\log p(x,y)$ $=\log\int_{z}\sum_{c=1}^{K}q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x)\,\frac{p_{\bm{\theta}}(x\mid z,c)\,p(y\mid c)\,p(z)\,p(c)}{q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x)}\,dz$
$\geq\sum_{c=1}^{K}q_{\bm{\phi}}(c\mid x)\int_{z}q_{\bm{\phi}}(z\mid x)\log\left(\frac{p_{\bm{\theta}}(x\mid z,c)\,p(y\mid c)\,p(z)\,p(c)}{q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x)}\right)dz.$
This gives the labelled ELBO:
$\mathcal{L}_{\text{label}}(x,y;\bm{\theta},\bm{\phi})$ $=\sum_{c=1}^{K}q_{\bm{\phi}}(c\mid x)\,\mathbb{E}_{q_{\bm{\phi}}(z\mid x)}\left[\log p_{\bm{\theta}}(x\mid z,c)\right]$
$\quad+\sum_{c=1}^{K}q_{\bm{\phi}}(c\mid x)\,\log p(y\mid c)$
$\quad-\mathrm{KL}(q_{\bm{\phi}}(z\mid x)\parallel p(z))-\mathrm{KL}(q_{\bm{\phi}}(c\mid x)\parallel p(c)).$
Supervision Term.
The term
$\sum_{c=1}^{K}q_{\bm{\phi}}(c\mid x)\,\log p(y\mid c)$
encourages the variational posterior over classes $q_{\bm{\phi}}(c\mid x)$ to concentrate on the true label $y$.
We typically define $p(y\mid c)$ as a deterministic distribution:
$p(y\mid c)=\delta_{c=y},$
where $\delta_{c=y}=1$ if $c=y$, and 0 otherwise. Consequently,
$\log p(y\mid c)=\begin{cases}0&\text{if }c=y,\\ -\infty&\text{otherwise}.\end{cases}$
Thus, if $q_{\bm{\phi}}(c\mid x)$ assigns non-zero probability to any class $c\neq y$, the corresponding term contributes $q_{\bm{\phi}}(c\mid x)(-\infty)$, which collapses the ELBO to -$\infty$. Maximization of the ELBO therefore forces the model to place almost all mass on the correct class. Formally, this reduces to:
$\sum_{c=1}^{K}q_{\bm{\phi}}(c\mid x)\,\log p(y\mid c)=\log q_{\bm{\phi}}(c=y\mid x).$
This is exactly the log-likelihood of the correct class under the variational posterior, equivalent to the negative cross-entropy loss used in supervised classification.
Interpretation.
This derivation shows that the supervision term not only promotes the correct label but also heavily penalises probability mass on incorrect labels. In doing so, it explicitly supervises the inference network $q_{\bm{\phi}}(c\mid x)$ to predict the semantic label $y$, rather than learning $c$ through unsupervised clustering, which may not align with actual class semantics.
Full Labelled ELBO.
[…]
Here, term (B) corresponds exactly to the supervised cross-entropy objective, with incorrect class assignments effectively penalized with infinite cost in the ELBO formulation.
A.2 Unlabelled Case (only $x$ known)
When $y$ is unobserved, the marginal log-likelihood becomes:
$\log p(x)=\log\int_{z}\sum_{c=1}^{K}p_{\bm{\theta}}(x\mid z,c)\,p(z)\,p(c)\,dz.$
Again, applying variational inference:
$\log p(x)$ $=\log\int_{z}\sum_{c=1}^{K}q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x)\,\frac{p_{\bm{\theta}}(x\mid z,c)\,p(z)\,p(c)}{q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x)}\,dz$
$\geq\sum_{c=1}^{K}q_{\bm{\phi}}(c\mid x)\int_{z}q_{\bm{\phi}}(z\mid x)\log\left(\frac{p_{\bm{\theta}}(x\mid z,c)\,p(z)\,p(c)}{q_{\bm{\phi}}(z\mid x)\,q_{\bm{\phi}}(c\mid x)}\right)dz.$
This gives the ELBO for unlabelled data:
$\mathcal{L}_{\text{unlabel}}(x;\bm{\theta},\bm{\phi})$ $=\sum_{c=1}^{K}q_{\bm{\phi}}(c\mid x)\,\mathbb{E}_{q_{\bm{\phi}}(z\mid x)}\left[\log p_{\bm{\theta}}(x\mid z,c)\right]$
$\quad-\mathrm{KL}(q_{\bm{\phi}}(z\mid x)\parallel p(z))-\mathrm{KL}(q_{\bm{\phi}}(c\mid x)\parallel p(c)).$
A.3 Total Objective
The full training objective over a labelled dataset $\mathcal{D}_{\ell}=\{(x_{i},y_{i})\}$ and unlabelled dataset $\mathcal{D}_{u}=\{x_{j}\}$ is:
$\mathcal{L}_{\text{total}}(\bm{\theta},\bm{\phi})=\sum_{(x,y)\in\mathcal{D}_{\ell}}\mathcal{L}_{\text{label}}(x,y;\bm{\theta},\bm{\phi})+\sum_{x\in\mathcal{D}_{u}}\mathcal{L}_{\text{unlabel}}(x;\bm{\theta},\bm{\phi}).$
@@ Appendix D Training hyper-parameters
Reconstruction loss weight 22
label loss weight 55
Gradient clipping (Max norm) 11
@@ Appendix F Imagination using GAN
Here, $z_{\text{reversed}}=\text{GRL}(E(x^{(l)}))$, and the gradient is scaled by $\lambda_{\text{dis}}$. The total Encoder-Decoder loss is: